OpenStudy (anonymous):

Suppose that a small pump can empty a swimming pool in 90 hours and that a large pump can empty a pool in 60 hours

OpenStudy (nurali):

\[\frac{ x }{ 60 }+\frac{ x }{ 90 }=1\] \[\frac{ 90x+60x }{ 5400 }=1\] \[150x=5400\] \[x=36\]

OpenStudy (anonymous):

Thank you :D

OpenStudy (nurali):

My Pleasure.